# A Survey of Text Watermarking in the Era of Large Language Models

### v3

@@ 5.3. Robustness
In the context of text watermarking, a crucial evaluation metric is its robustness against watermark removal attacks. A watermark removal attack refers to the process of altering watermarked text in an attempt to erase the embedded watermark. If a watermarked text still has a high probability of being detected following a Watermark Removal Attack, then the text watermarking algorithm is considered highly robust.
Current assumptions for watermark removal attacks are based on black-box access to the watermarking algorithm, meaning the method of watermark generation is unknown and there is no access to the watermark detector during the attack. This is primarily because, under white-box access, potent watermark removal attack algorithms can be easily developed to remove most watermarks, as mentioned by (Kirchenbauer et al., 2023b) in their anti-watermark schema. This represents a limitation of all current watermarking algorithms. Although it is possible that watermarks may not withstand watermark removal attacks under white-box access, exploring these possibilities or designing robust watermark algorithms under white-box settings remains a direction for future research.
Under the premise of black-box access, numerous watermark removal attacks have been developed to erase the watermark by modifying watermarked text. Based on the granularity of textual modifications, we categorize these methods into character-level attacks, word-level attacks, and document-level attacks. In the following part of this section, we will discuss the implementation of these attacks and the robustness of current text watermarking methods against them.
@@
Modifying characters in text without altering any actual words is a relatively straightforward method of watermark removal attack. One approach involves directly perturbing certain characters to create spelling errors. However, this method is easily detectable and can compromise the quality of the text. An alternative strategy involves replacing characters with visually similar Unicode IDs, leveraging the fact that many Unicode IDs correspond to identical or visually indistinguishable characters. This technique is also known as a Homoglyph attack (Gabrilovich and Gontmakher, 2002). Although such methods are difficult to detect by human observation, they can still be mitigated by various canonicalization techniques. Therefore, normalization preprocessing before passing through watermark detectors is crucial.
@@
The advantage of a character-level attack is its simplicity and effectiveness. However, its drawback is that it is easily detectable or can be eliminated by some simply designed algorithms. Therefore, it is not a reliable watermark removal attack in all scenarios.
@@
Regarding watermarking during logits generaion methods (section 4.2), the basic attack effect is that the watermarked tokens (e.g. green tokens in Kirchenbauer et al. (2023a)) in the text will be disturbed. For zero-bit algorithms (Kirchenbauer et al., 2023a; Zhao et al., 2023a; Liu et al., 2023b, c), insertion or replacement of non-watermarked tokens as well as deletion of watermarked tokens result in decreased proportion of watermarked token in the detected text. Consequently, the detector whose result relies on calculating the ratio of watermarked tokens will output lower detection scores, causing more false negative cases (Christ et al., 2023; Kuditipudi et al., 2023; Kirchenbauer et al., 2023a; Zhao et al., 2023a). For multi-bit algorithms (Wang et al., 2023b; Yoo et al., 2023a), disturbance of the tokens in a text part originally embedded with a message will result in wrongly decoded message (Wang et al., 2023b; Yoo et al., 2023a; Abdelnabi and Fritz, 2021; Yoo et al., 2023b). That is because the decoding process is to choose the message with the highest probability among all the candidates. In that case, altering one or more words in a text fragment could cause the decoding probability to shift to an unintended one. For algorithms like (Kirchenbauer et al., 2023a; Zhao et al., 2023a; Wang et al., 2023b), there is extra effect because these algorithms rely on preceding tokens to embed watermark and decode watermark. Even though an original watermarked token survived all three editing attacks, the changed surrounding would cause detection computation problems on that token, and eventually lead to failed watermark decoding on such tokens (Kirchenbauer et al., 2023a).
Usually, to achieve significant performance drop, sustainable word-level attacks on the text is required. That is, the attack rate should be large enough. For example, to decrease detection F1 score below 80%, a minimal attack rate of 30% is required, which requires performing deletion or replacement every 3 words (Yang et al., 2023). The main problems brought by a large attack rate are the totally different semantics and noticeably degraded text quality (Abdelnabi and Fritz, 2021; Kirchenbauer et al., 2023a; Yang et al., 2023). Word deletion with high attack rate may even leave each sentence incomplete. Often, the text under such heavy word-level attacks is not ideal in real world, where the watermark-removed text should still be usable and understandable.
@@
Word-level Attack during Text Generation. Due to the inevitable impact on text quality during word-level attack to existing texts, particularly when these modifications are extensive, recent work has begun to explore word-level attacks during the text generation phase. These methods primarily targets on watermarking for LLMs. A notable example is the emoji attack (Goodside, 2023), which prompts the LLM to generate emojis between each token and then remove the emojis after generation. This attack is effective when watermark embedding process depends on the preceding token. After the emojis are removed, detector would fail to recognize the watermarked tokens as emojis are supposed to carry the watermarking information of a subsequent token (Kirchenbauer et al., 2023a) (Sun et al., 2023). For example, a user could request the LLM to generate "Λ" between every word, resulting in a sentence looking like "There Λ are Λ some Λ apples Λ here". Assuming the word "apples" is watermarked, and the proper detection computation of this word needs its prefix, the emoji "Λ". However, the user will remove the emojis before using this machine-generated text, making the prefix of "apples" become the word "some". Detection using a different prefix will be inaccurate, and such miscalculation is applied in every word of the watermarked text.
The advantage of such attacks is their near-complete erasure of certain watermarking methods (Kirchenbauer et al., 2023a; Christ et al., 2023). However, they have two main disadvantages. Firstly, they depend heavily on the strong adherence ability of Large Language Models to follow instructions. Powerful LLMs like ChatGPT and Claude can effectively execute emoji attacks, ensuring quality text output, but less capable LLMs may fail to follow instructions, leading to unreasonable text generation. Kirchenbauer et al. (2023a) proposed a solution to include such prompts during LLM fine-tuning so that LLM could be taught to reject such requests from users. However, the efficacy of such fine-tuning is not guaranteed and may potentially diminish the capabilities of Large Language Models. Secondly, the success of these attack methods is strongly linked to the watermark generation process. For instance, the emoji attack assumes that the current token’s watermark status (whether it belongs to the "green list) is dependent on the hash value of the previous token. However, methods like those proposed by (Liu et al., 2023c), which do not rely on the preceding token’s hash value but use the embedding of generated text to determine the watermark status, are minimally affected by the emoji attack.
@@
Although word-level attack could modify individual words to alter the remove the text watermark, their scope and depth of impact are relatively limited. In contrast, document-level attack employ a more comprehensive and profound approach. These attacks go beyond mere word modifications in a document, encompassing broader adjustments to the entire content and structure of the document. Common document-level attacks include semantic-preserving rewrites, also known as rewrite attacks, and the insertion of watermark text into existing human text, termed copy-paste attacks.
@@
Rewrite Attack. The rewrite attack offers comprehensive and in-depth modifications to text, yet its implementation is more challenging compared to the word-level approach. Early methods of rewrite attacks often employed a re-translation strategy (Yang et al., 2023), which involves translating the text into another language and then back into the original language. This method exploits subtle changes that may occur during the translation process to achieve text rewriting. While the re-translation method was widely used initially, it has a drawback: the double translation can introduce additional errors, leading to significant semantic deviations from the original text. Consequently, specialized language models for text rewriting, such as Dipper (Krishna et al., 2023), have been developed. These models provide higher quality text and allow for configurable degrees of modification during the rewriting process. With the recent popularity of ChatGPT, many studies and practices have started using it for text rewriting, most commonly the gpt-3.5-Turbo version. Its advantage lies in requiring only a specific prompt for the text rewrite attack, such as "please rewrite the following text:". For more realistic scenario validation, manual rewriting is also an option. Manual rewriting offers more precise semantic preservation and more natural language expression but has the main disadvantage of being costlier, especially when handling large amount of text.
The effectiveness of rewrite attacks varies for different types of watermarking methods. For format-based watermarking approaches (Brassil et al., 1995; Sato et al., 2023; Por et al., 2012; Rizzo et al., 2016), rewrite attack is disastrous as any homoglyph will be replaced to the standard language used by the LLM or human in the output. For other algorithms that embed watermarks based on text content, the impact of rewrite attacks is uncertain and may depend on three factors: whether the watermark detection is related to the sequence of tokens, the strength of watermark implantation (i.e., the extent of text modification), and the length of the text. Since rewrite attack could easily disrupt the order of tokens but struggles to replace all tokens, watermark detection method that do not rely on token sequences, like the method proposed by Zhao et al. (2023a), tend to be more robust. Conversely, methods like those of Kuditipudi et al. (2023), where robustness is affected by token sequence, show different levels of susceptibility. Additionally, if the watermark text requires significant modifications from the unwatermarked text, it enhances the robustness against rewrite attacks. The length of the watermarked text is also crucial. Experiments by Kirchenbauer et al. (2023a) show that watermarked texts exceeding 600 tokens are generally robust against all rewrite attacks. It is noteworthy that for humans, erasing watermarks from texts ranging from 400 to 800 tokens through rewriting is exceptionally challenging.
@@
There are some designs to further increase algorithm robustness against rewrite attack as well. 1) Decrease the token-level dependency. The dependency of watermark algorithms on specific contexts can compromise their effectiveness when the original text is rewritten, as the original context may not remain intact. Take the watermarking during logits generation methods as example, if the partitioning of a token’s red-green list relies on $m$ adjacent words, the detection rate significantly decreases after an attack when $m$ is not very small. Hence, watermark algorithms should reduce reliance on neighboring words. This can be achieved by decreasing the window size $m$ (Kirchenbauer et al., 2023a), or by using hashing schemes that only involve a subset of words within the window (Kirchenbauer et al., 2023b). Some studies have even reduced the window size $m$ to zero (Zhao et al., 2023a). 2) Replace with semantic-level dependency. A better approach is for the watermark algorithm to refer to the general semantics around the watermarked token (Liu et al., 2023c; Yoo et al., 2023b) instead of the exactly identical tokens or words. That is to say, for a group of altered but semantically similar contexts, the watermark generator and detector would produce the same result. Unless the semantics is drastically altered, the watermark detection could maintain a higher accuracy against rewrite attacks. One can even train the watermark generator to produce closer results on semantically close texts ; 3) Increase embedding space for each message. The paraphrasing attack would change the appearance of a watermarked text fragment and disable accurate detection in that fragment. If the message is only embedded in one short text piece, the message would be easily erased if the re-writer considered that part redundant. Thus, increasing the number or length of text fragments that are embedded with the same message would intuitively increase the success extraction rate (Munyer and Zhong, 2023; Wang et al., 2023b).
@@
It is worth noting that although human writers are stronger paraphrasers than machines. However, there are significant differences in individuals’ ability to rewrite text, and moreover, when the text is lengthy, humans are often powerless.
Copy-paste Attack. Copy-paste attack is to surround the target text with distraction text, which in this context equal to the watermarked and non-watermarked text respectively. Such attack will result in a much longer text as compared to the original watermarked text. This attack aims to test if the low ration setting can cause algorithm effectiveness to drop.
@@
The copy-paste attack primarily diminishes the detectability of watermarks in text by diluting their concentration. The efficacy of this attack is significantly influenced by the proportion of watermarked text. For instance, when the watermarked text constitutes a mere 10% of the total, the attack’s impact tends to surpass that of most rewriting attacks (Kirchenbauer et al., 2023b). However, as the share of watermarked text increases to, say, 25%, its effectiveness in undermining certain watermarking algorithms (Kirchenbauer et al., 2023a) becomes comparable to that of some rewriting attacks. Similar to rewriting attacks, lengthening the text can enhance the reliability of watermark detection, particularly in the context of copy-paste attacks.
@@
However, copy-paste attacks may be identified by certain specific watermark detection methods. For example, Kirchenbauer et al. (2023b) mentioned a windowed test to calculate the level of watermarking in a region of the text instead of the whole text. The idea is to detect watermarked text that is inserted into an existing text. This should be effective for copy-paste attack.
@@ 5.4. Unforgeability
The discussion on watermark unforgeability could be differentiated into two distinct scenarios: private detection and public detection. In the private detection scenario, the watermark detection method is kept confidential, accessible only to specific users or institutions, or indirectly provided to users via an API (Kirchenbauer et al., 2023a). This setup’s advantage lies in increasing the difficulty for attackers to obtain and analyze the detection algorithm, as they lack direct access to the detection mechanism. Conversely, in the public detection scenario, the watermark detection algorithm is publicly available, allowing anyone to access and utilize these algorithms. The benefit of this approach is its transparency and ease of integration, making it widely applicable in various contexts where authenticity verification is needed. However, this also implies that attackers can more easily study the detection algorithms and seek methods to breach the watermark. Therefore, the requirements for unforgeability in public detection scenarios might be higher.
@@ 5.4.1. Unforgeability in Privately Detectable Scenario
In private detection scenarios, the imperceptibility of text watermarking algorithms is crucial for ensuring unforgeability due to the limited detection capabilities of users. Imperceptibility refers to the watermark’s impact on the original content being nearly undetectable in its statistical distribution, which is essential for ensuring the watermark’s non-forgery. If users are unaware of the watermark in the text, they cannot forge it. Therefore, some studies have focused on testing unforgeability. For instance, Abdelnabi and Fritz (2021) collected texts with and without watermarks and trained classifiers to try to distinguish between these two categories. The purpose of this test was to see if the classifiers could effectively identify which texts contained watermarks. The performance of the classifiers serves as a measure of the watermarking algorithm’s imperceptibility. Ideally, classifiers should find it challenging to differentiate between watermarked and non-watermarked texts. If the classifier cannot accurately differentiate, this indicates that the watermarking algorithm performs well in terms of imperceptibility, thereby enhancing its unforgeability.
@@
When text watermarking algorithms fail to achieve complete imperceptibility, or when attackers suspect or infer the presence of watermarks in the text, they may resort to statistical methods to extract the watermarking algorithm’s generation methods. These statistical attack methods primarily rely on the analysis of statistical traces or patterns that the watermarking algorithm might leave in the text, which generally requires sufficient prior knowledge about the method of watermark insertion. For instance, Sadasivan et al. (2023) proposed a spoofing attack algorithm that can target numerous watermarking during logits generation methods (Kirchenbauer et al., 2023a; Zhao et al., 2023a). Specifically, this attack involves calculating the frequency of tokens in a text under a fixed prefix. In watermarked texts, the frequency of these tokens may differ from normal texts. For instance, by analyzing the frequency of tokens following ‘the’, tokens that appear more frequently might be identified as part of the watermark (green list (Kirchenbauer et al., 2023a)). The effectiveness of this attack is closely related to the complexity of the watermarking method. If a watermarking algorithm operates in a complex manner on a text, statistical-based attacks become less effective. This suggests that increasing the complexity of watermarking algorithms is key to preventing such attacks. For example, Liu et al. (2023c) proposed using text semantic information associated with the watermarking rule, which effectively resists spoofing attacks.
In the context of private watermark detection scenarios, there are measures that can enhance the unforgeability of watermark algorithms. Firstly, it is necessary to limit the detection frequency. When providing watermark detection services via an API, mechanisms can be designed to restrict this frequency. This helps prevent attackers from understanding and analyzing the watermark algorithm through extensive trial and error, thereby reducing the effectiveness of the previously mentioned spoofing attacks (Sadasivan et al., 2023). Secondly, bolstering network security is crucial for preventing the theft of watermark rules. Protecting the API and backend systems from hacker intrusions includes, but is not limited to, using encrypted communications, regularly updating security vulnerabilities, and implementing intrusion detection systems. Thirdly, guarding against social engineering attacks is essential. Social engineering attacks often manipulate people’s trust or induce them to disclose information. Establishing strict internal security protocols and verification processes is vital to prevent unauthorized information (e.g. the key of watermark) disclosure.

### v4

@@ 5.3. Robustness
Another key metric for evaluation is the robustness against watermark removal attacks. These attacks entail modifying watermarked text to erase the embedded watermark. A watermarking algorithm is considered highly robust if the watermark remains detectable post-attack.
Presently, watermark removal attack models assume black-box access to the watermarking algorithm. Here, the watermark generation method is undisclosed, and there is no access to the watermark detector during the attack. This assumption is due to the ease of developing effective removal strategies under white-box access, which could potentially eliminate most watermarks, as highlighted by Kirchenbauer et al. (2023b) in their anti-watermark framework. In scenarios with black-box access, several watermark removal attacks have been developed with the aim of erasing the watermark through textual modifications. These methods are categorized based on modification granularity, including character-level, word-level, and document-level attacks. The subsequent sections will explore these attack implementations and evaluate the resilience of text watermarking techniques against them.
@@
Character modification in text, without altering words, is a basic strategy for watermark removal attacks. One method involves introducing spelling errors by perturbing characters, but this can be easily identified and may degrade text quality. An alternative strategy involves replacing characters with visually similar Unicode IDs, leveraging the fact that many Unicode IDs correspond to identical or visually indistinguishable characters. This technique is also known as a Homoglyph attack (Gabrilovich and Gontmakher, 2002). Although such methods are difficult to detect by human observation, they can still be mitigated by various canonicalization techniques (Kirchenbauer et al., 2023a). Therefore, normalization preprocessing before passing through watermark detectors is crucial.
@@
A character-level attack offers simplicity and effectiveness as its advantage. However, its drawback is that it is easily detectable or can be eliminated by some simply designed algorithms. Therefore, it is not a reliable watermark removal attack in all scenarios.
@@
Regarding watermarking during logits generaion methods (section 4.2), the basic attack effect is that the watermarked tokens (e.g. green tokens in Kirchenbauer et al. (2023a)) in the text will be disturbed. Watermark methods (Kirchenbauer et al., 2023a; Zhao et al., 2023a; Wang et al., 2023b) that rely on preceding tokens to determine the current token’s watermark status are more vulnerable to word-level attacks. This is because word-level attacks not only change the current token but also modify its preceding tokens, thereby altering the detection status of the current token.
Usually, to achieve significant performance drop, sustainable word-level attacks on the text is required. That is, the attack rate should be large enough. For example, to decrease detection F1 score below 80%, a minimal attack rate of 30% is required, which requires performing deletion or replacement every 3 words (Yang et al., 2023). High attack rates drastically change semantics and degrade text quality (Abdelnabi and Fritz, 2021; Kirchenbauer et al., 2023a; Yang et al., 2023). Word deletion with high attack rate may even leave each sentence incomplete. Often, the text under such heavy word-level attacks is not ideal in real world, where the watermark-removed text should still be usable and understandable.
@@
Word-level Attack during Text Generation. Due to the inevitable impact on text quality during word-level attack to existing texts, particularly when these modifications are extensive, recent work has begun to explore word-level attacks during text generation. These methods target watermarking algorithms for LLMs. A notable example is the emoji attack (Goodside, 2023), which prompts the LLM to generate emojis between each token and then remove the emojis after generation.
This attack is effective when watermark embedding process depends on the preceding token. Once the emojis are removed, the detector would fail to recognize the watermarked tokens, as emojis are intended to convey the watermarking information of a subsequent token (Kirchenbauer et al., 2023a; Sun et al., 2023). For example, a user could request the LLM to generate "Λ" between every word, resulting in a sentence looking like "There Λ are Λ some Λ apples Λ here". Assuming the word "apples" is watermarked, and the proper detection computation of this word needs its prefix, the emoji "Λ". However, the user will remove the emojis before using this machine-generated text, making the prefix of "apples" become the word "some". Detection using a different prefix will be inaccurate, and such miscalculation is applied in every word of the watermarked text.
The advantage of such attacks is their near-complete erasure of certain watermarking methods (Kirchenbauer et al., 2023a; Christ et al., 2023). However, they have two main disadvantages. First, their effectiveness hinges on the LLM’s ability to adhere to directives. Advanced LLMs like ChatGPT and Claude can successfully perform emoji attacks, ensuring coherent text generation, but less advanced models might struggle, resulting in illogical outputs. Second, the success of these attacks is intricately linked to the watermark generation mechanism. For instance, the emoji attack presupposes that a token’s watermark status depends on the preceding token’s hash value. However, alternative approaches, like those by (Liu et al., 2023c), which utilize generated text embeddings instead of relying on preceding token hashes, remain largely unaffected by emoji attacks.
@@
While word-level attacks alter or remove text watermarks through individual word modifications, their impact is comparatively confined. Conversely, document-level attacks adopt a holistic and in-depth strategy. They extend beyond mere word alterations in a document, encompassing extensive modifications to both content and structure. Typical document-level attacks involve semantic-preserving rewrites, often referred to as rewrite attacks, and the integration of watermark text into pre-existing human text, known as copy-paste attacks.
@@
Rewrite Attack. The rewrite attack offers comprehensive and in-depth modifications to text, yet its implementation is more challenging compared to the word-level approach. Early methods of rewrite attacks often employed a re-translation strategy (Yang et al., 2023), which involves translating the text into another language and then back into the original language. This method exploits subtle changes that may occur during the translation process to achieve text rewriting. While the re-translation method was widely used initially, its limitation lies in the potential introduction of additional errors during double translation, which can result in significant semantic deviations from the original text. Consequently, specialized language models for text rewriting, such as Dipper (Krishna et al., 2023), have been developed. These models provide higher quality text and allow for configurable degrees of modification during the rewriting process. With the recent popularity of ChatGPT, many studies and practices have started using it for text rewriting, most commonly the gpt-3.5-Turbo version. Its advantage lies in requiring only a specific prompt for the text rewrite attack, such as "please rewrite the following text:". For more realistic scenario validation, manual rewriting is also an option. Manual rewriting offers more precise semantic preservation and more natural language expression but has the main disadvantage of being costlier, especially when handling large amount of text.
The effectiveness of rewrite attacks varies for different types of watermarking methods. Format-based approaches (Brassil et al., 1995; Sato et al., 2023; Por et al., 2012; Rizzo et al., 2016) are particularly vulnerable, as homoglyphs are often replaced with standard language tokens by LLMs or humans in the output. For other algorithms that embed watermarks based on text content, the impact of rewrite attacks depends on three factors: token sequence dependence, watermark strength, and text length.
Firstly, watermark detection methods that do not rely on token order, like Zhao et al. (2023a)’s approach, demonstrate greater robustness since rewrite attacks can disrupt token sequences but are less effective at replacing all tokens. In contrast, algorithms like Kuditipudi et al. (2023)’s, where robustness hinges on token sequences, exhibit varying susceptibility levels to such attacks. Secondly, the watermark strength, or the extent of text modification, is crucial. Watermarking that necessitates substantial alterations from the original text tends to enhance resistance against rewrite attacks. Lastly, the length of the watermarked text is a significant factor. Empirical evidence from Kirchenbauer et al. (2023a) indicates that texts with over 600 tokens generally maintain robustness against rewrite attacks. It is also observed that manually erasing watermarks from texts of 400 to 800 tokens through rewriting poses a considerable challenge for humans.
@@
There are some designs to further increase algorithm robustness against rewrite attack as well. 1) Decrease the token-level dependency. The dependency of watermark algorithms on specific contexts can compromise their effectiveness when the original text is rewritten, as the original context may not remain intact. Take the watermarking during logits generation methods as example, if the partitioning of a token’s red-green list relies on $m$ adjacent words, the detection rate significantly decreases after an attack when $m$ is not very small. Hence, watermark algorithms should reduce reliance on neighboring words. This can be achieved by decreasing the window size $m$ (Kirchenbauer et al., 2023a), or by using hashing schemes that only involve a subset of words within the window (Kirchenbauer et al., 2023b). Some studies have even reduced the window size $m$ to zero (Zhao et al., 2023a). 2) Replace with semantic-level dependency. A better approach is for the watermark algorithm to refer to the general semantics around the watermarked token (Liu et al., 2023c; Yoo et al., 2023b) instead of the exactly identical tokens or words. That is to say, for a group of altered but semantically similar contexts, the watermark generator and detector would produce the same result. Unless the semantics is drastically altered, the watermark detection could maintain a higher accuracy against rewrite attacks. One can even train the watermark generator to produce closer results on semantically close texts ;
@@
It is worth noting that although human writers are stronger paraphrasers than machines, there are significant differences in individuals’ ability to rewrite text. Moreover, when the text is lengthy, humans are often powerless.
@@
Copy-paste Attack. Copy-paste attack is to surround the target text (watermarked text) with distraction text (non-watermarked text). This attack aims to test whether a low ratio of watermarked text can cause a drop in algorithm effectiveness. The copy-paste attack primarily diminishes the detectability of watermarks in text by diluting their concentration. The efficacy of this attack is significantly influenced by the proportion of watermarked text. For instance, when the watermarked text constitutes a mere 10% of the total, the attack’s impact tends to surpass that of most rewriting attacks (Kirchenbauer et al., 2023b). However, as the share of watermarked text increases to, say, 25%, its effectiveness in undermining certain watermarking algorithms (Kirchenbauer et al., 2023a) becomes comparable to that of some rewriting attacks. Similar to rewriting attacks, lengthening the text can enhance the reliability of watermark detection, particularly in the context of copy-paste attacks.
@@
However, copy-paste attacks may be identified by certain specific watermark detection methods. For example, Kirchenbauer et al. (2023b) mentioned a windowed test to calculate the level of watermarking in a region of the text instead of the whole text. It is specifically designed to effectively detect watermarked text that is inserted into existing text, making it applicable for addressing copy-paste attacks.
@@ 5.4. Unforgeability
The discussion on watermark unforgeability could be differentiated into two distinct scenarios: private detection and public detection. In the private detection scenario, the watermark detection method is kept confidential, accessible only to specific users or institutions, or indirectly provided to users via an API (Kirchenbauer et al., 2023a). This setup’s advantage lies in increasing the difficulty for attackers to obtain and analyze the detection algorithm, as they lack direct access to the detection mechanism. Conversely, in the public detection scenario, the watermark detection algorithm is publicly available, allowing anyone to access and utilize these algorithms. In this scenario, attackers can more easily study the detection algorithms and seek methods to breach the watermark. Therefore, the requirements for unforgeability in public detection scenarios might be higher.
@@ 5.4.1. Unforgeability in Privately Detectable Scenario
In private detection scenarios, the imperceptibility of text watermarking algorithms is crucial for ensuring unforgeability due to the limited detection capabilities of users. Imperceptibility refers to the watermark’s impact on the original content being nearly undetectable in its statistical distribution, which is essential for ensuring the watermark’s non-forgery. If users are unaware of the watermark in the text, they cannot forge it. Therefore, some studies have focused on testing imperceptibility. For instance, Abdelnabi and Fritz (2021) collected texts with and without watermarks and trained classifiers to try to distinguish between these two categories. The purpose of this test was to see if the classifiers could effectively identify which texts contained watermarks. The performance of the classifiers serves as a measure of the watermarking algorithm’s imperceptibility. The classifiers’ efficacy in detecting watermarks serves as a proxy for the algorithm’s imperceptibility, with ideal outcomes being their inability to distinguish between the two categories.
@@
When text watermarking algorithms fail to achieve complete imperceptibility, or when attackers suspect or infer the presence of watermarks in the text, they may resort to statistical methods to extract the details of the watermark generator. These statistical attack methods primarily rely on the analysis of statistical traces or patterns that the watermarking algorithm might leave in the text, which generally requires sufficient prior knowledge about the method of watermark insertion. For instance, Sadasivan et al. (2023) proposed a spoofing attack algorithm that can target numerous watermarking during logits generation methods (Kirchenbauer et al., 2023a; Zhao et al., 2023a). Specifically, this attack involves calculating the frequency of tokens in a text under a fixed prefix. In watermarked texts, the frequency of these tokens may differ from normal texts. For instance, by analyzing the frequency of tokens following ‘the’, tokens that appear more frequently might be identified as part of the watermark (green list (Kirchenbauer et al., 2023a)). The effectiveness of this attack is closely related to the complexity of the watermarking method. If a watermarking algorithm operates in a complex manner on a text, statistical-based attacks become less effective. This suggests that increasing the complexity of watermarking algorithms is key to preventing such attacks. For example, Liu et al. (2023c) proposed using text semantic information associated with the watermarking rule, which effectively resists spoofing attacks.
In the context of private watermark detection scenarios, several strategies are imperative. First, detection frequency must be limited. For watermark detection services via APIs, frequency-restricting mechanisms can reduce attackers’ capacity to analyze the algorithm through repeated attempts, thus mitigating spoofing attacks (Sadasivan et al., 2023). Second, enhancing network security is critical to safeguard watermark rules against theft. This includes encrypted communications, regular security updates, and intrusion detection systems. Third, defending against social engineering attacks is essential. These attacks exploit trust or coerce information disclosure. Implementing stringent internal security protocols and verification processes is crucial to prevent unauthorized release of sensitive data, such as watermark keys.